OpenStudy (anonymous):

2 sqrt 5 + 3 sqrt 7^2

OpenStudy (diyadiya):

is this your question ? \[2\sqrt{5}+3\sqrt{7^2}\]

OpenStudy (anonymous):

let me rewrite (2 sqrt 5 + 3 sqrt 7)^2

OpenStudy (diyadiya):

\[(2\sqrt{5}+3\sqrt{7})^2\] Exapand using the formula (a+b)^2

OpenStudy (anonymous):

>.< erg i hate math it tells me to justify each step >.<

OpenStudy (diyadiya):

\[(a+b)^2=(a)^2+2(a)(b)+(b)^2\] can you substitute a and b here \[a=2\sqrt{5} \ \ b=3\sqrt{7}\]

OpenStudy (diyadiya):

\[(2\sqrt{5})^2+2 \times 2\sqrt{5} \times 3\sqrt{7}+(3\sqrt{7})^2\]

OpenStudy (diyadiya):

\[(2\sqrt{5}+3\sqrt{7})^2\] this is of the form \[(a+b)^2\]Did you understand this?

OpenStudy (anonymous):

yes

OpenStudy (diyadiya):

Here \[a=2\sqrt{5} \ \ b=3\sqrt{7}\]

OpenStudy (diyadiya):

understood?

OpenStudy (anonymous):

yea i got it thanks

OpenStudy (diyadiya):

sure?

OpenStudy (anonymous):

yep thanks so much

OpenStudy (diyadiya):

Welcome